OpenStudy (anonymous):

What is the y-intercept of the equation 2x + 3y = 12? -1/2 4 6 12

OpenStudy (anonymous):

its 4 get the equation in the form y=mx + c and

jimthompson5910 (jim_thompson5910):

To find the y-intercept, plug in x = 0 and solve for y 2x + 3y = 12 2*0 + 3y = 12 0 + 3y = 12 3y = 12 y = 12/3 y = 4 So the y-intercept as a point is (0,4)

OpenStudy (anonymous):

c is the intercept

OpenStudy (anonymous):

recast it as y = m x + b, and b is your answer.